# Semi-Bandit Learning for Monotone Stochastic Optimization

### v1

@@ 3 Applications
Proof.
We use the following result on $b$-transportations in the proof of this lemma.
$b$-Transportation. Given a bipartite graph $G=(V,E)$, and a vector $b\in\mathbb{R}^{V}_{+}$, a $b$-transportation is a vector $f\in\mathbb{R}^{E}_{+}$ with $f(\delta(v))=b_{v}$ for each $v\in V$. Here $f(\delta(v))$ denotes the sum of $f$-values of edges incident on $v$. The following theorem (due to [Rad43], see also Theorem 21.11 in [${}^{+}$03]) characterizes a $b$-transportation.
Theorem 3.2.
Let $G=(V,E)$ be a bipartite graph and let $b\in\mathbb{R}^{V}_{+}$. Then there exists a $b$-transportation if and only if $b(C)\geq\frac{1}{2}\cdot b(V)$ for every vertex cover $C$.
Consider an arbitrary instance $\mathcal{I}$ and product probability distributions $\mathbf{{D}}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ and $\mathbf{{E}}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ where $\mathbf{{D}}\preceq_{\mathsf{SD}}\mathbf{{E}}$. To prove the lemma, we need the following intermediate result.
@@
Coupling distributions under stochastic dominance. Consider r.v.s $\mathsf{X}$ and $\mathsf{Y}$ with finite support $[k]=\{1,\ldots,k\}$, and distributed according to $\mathcal{D}$ and $\mathcal{E}$ respectively. Furthermore, suppose that $\mathcal{D}\preceq_{\mathsf{SD}}\mathcal{E}$. Then, we will show that there exist values $\left\{f_{\ell_{1},\ell_{2}}:k\geq\ell_{1}\geq\ell_{2}\geq 1\right\}$ such that:
$\sum_{\ell_{2}:\ell_{2}\leq\ell_{1}}f_{\ell_{1},\ell_{2}}=\mathbf{P}_{\mathcal{E}}(\mathsf{Y}=\ell_{1}),\,\forall\ell_{1}\in[k]\quad\mbox{ and }\quad\sum_{\ell_{1}:\ell_{1}\geq\ell_{2}}f_{\ell_{1},\ell_{2}}=\mathbf{P}_{\mathcal{D}}(\mathsf{X}=\ell_{2}),\,\forall\ell_{2}\in[k].$
The $\{f_{\ell_{1},\ell_{2}}\}$ values allow us to couple the realizations of r.v.s $\mathsf{X}\sim\mathcal{D}$ and $\mathsf{Y}\sim\mathcal{E}$. We can obtain these values by constructing a $b$-transportation instance on an appropriately defined bipartite graph: consider the bipartite graph $G=(U_{E}\cup U_{D},E)$ where the vertex sets represent the support of the random variable under distributions $\mathcal{E}$ and $\mathcal{D}$ respectively. Specifically, let $U_{E}=\{e_{1},\ldots,e_{k}\}$ and $U_{D}=\{d_{1},\ldots,d_{k}\}$, and $E=\{(e_{\ell_{1}},d_{\ell_{2}}):\ell_{1}\geq\ell_{2}\}$. Consider a $b$-value of $\mathbf{P}_{\mathsf{Y}\leftarrow\mathcal{E}}(\mathsf{Y}=\ell)$ on $e_{\ell}$, and a $b$-value of $\mathbf{P}_{\mathsf{X}\leftarrow\mathcal{D}}(\mathsf{X}=\ell)$ on $d_{\ell}$ for all $\ell=1,\ldots,k$. To apply Theorem 3.2, it suffices to show that for any vertex cover $C$ of $G$, we have $b(C)\leq\frac{1}{2}\cdot b(V)=1$ (since the $b$-values correspond to probabilities). Let $C$ be a fixed but arbitrary vertex cover of $G$, and observe that if $e_{\ell}\notin C$, then $C$ must contain $\{d_{\widehat{\ell}}:\widehat{\ell}\leq\ell\}$ in order to cover all edges incident on $e_{\ell}$. Consequently, any minimal vertex cover must have the form $\{e_{k},e_{k-1},\ldots,e_{\overline{\ell}},d_{\overline{\ell}-1},\ldots d_{1}\}$ where ¯$\overline{\ell}\in\{1,\ldots,k+1\}$. Let $\overline{C}\subseteq C$ be a minimal vertex cover, which must have the form described above; that is, $\overline{C}=\{e_{k},e_{k-1},\ldots,e_{\overline{\ell}},d_{\overline{\ell}-1},\ldots d_{1}\}$ for some ¯$\overline{\ell}\in\{1,\ldots,k+1\}$. Then,
$b(C)\geq b(\overline{C})=\sum_{\ell_{1}\geq\overline{\ell}}b(e_{\ell_{1}})+\sum_{\overline{\ell}>\ell_{2}\geq 1}b(d_{\ell_{2}})\geq\sum_{\ell_{1}\geq\overline{\ell}}b(d_{\ell_{1}})+\sum_{\overline{\ell}>\ell_{2}\geq 1}b(d_{\ell_{2}})=1.$
The last inequality follows from stochastic dominance $\mathcal{D}\preceq_{\mathsf{SD}}\mathcal{E}$ because:
$\sum_{\ell_{1}\geq\overline{\ell}}b(e_{\ell_{1}})=\mathbf{P}_{\mathsf{Y}\leftarrow\mathcal{E}}(\mathsf{Y}\geq\overline{\ell})\geq\mathbf{P}_{\mathsf{X}\leftarrow\mathcal{D}}(\mathsf{X}\geq\overline{\ell})=\sum_{\ell_{1}\geq\overline{\ell}}b(d_{\ell_{1}}).$
Thus, the condition of Theorem 3.2 holds, and so there exists a $b$-transportation $f:E\to\mathbf{R}$ for $G$. See Figure 1 for an example.
Figure 1: Suppose r.v.s $\mathsf{X}$ and $\mathsf{Y}$ have support $\{1,\ldots,5\}$. Let $\mathsf{X}\sim\mathcal{D}$ where $\mathcal{D}=\{1:0.2,2:0.1,3:0.1,4:0.3,5:0.3\}$. Here, each entry denotes the probability of the r.v. taking on the corresponding value; that is, $\mathbf{P}(\mathsf{X}=2)=0.1$. Similarly, let $\mathsf{Y}\sim\mathcal{E}$ where =$\mathcal{E}=\{1:0,2:0.1,3:0.3,4:0.2,5:0.4\}$. It is easy to check that $\mathcal{D}\preceq_{\mathsf{SD}}\mathcal{E}$. The figure depicts the corresponding $b$-transportation instance, and values that couple distributions $\mathcal{D}$ and $\mathcal{E}$ (as per Lemma 3.1).
@@
We are now ready to complete the proof. For each item $i\in[n]$, let $f_{\ell_{1},\ell_{2}}^{i}$ denote coupled values for distributions $\mathcal{D}_{i}\preceq_{\mathsf{SD}}\mathcal{E}_{i}$, constructed as per the above procedure. Let $\sigma_{\mathbf{{E}}}$ denote the optimal policy for $\mathcal{I}$ under distribution $\mathbf{{E}}$, and let $\mathtt{OPT}_{\mathcal{I}}(\mathbf{{E}})$ denote its reward. Recall that a policy is given by a decision tree, where each node is labeled by an item to probe next, and the branches out of a node correspond to the random realization of the probed item. We construct a policy $\widehat{\sigma}$ that works under distribution $\mathbf{{D}}$, by closely following $\sigma_{\mathbf{{E}}}$. When policy $\sigma_{\mathbf{{E}}}$ probes some item $i$, policy $\widehat{\sigma}$ does the same. Now, suppose that $\widehat{\sigma}$ observes $C_{i}=\ell_{2}\in[k]$. Then, for each $\ell_{1}\geq\ell_{2}$, with probability $\frac{f_{\ell_{1},\ell_{2}}^{i}}{\mathbf{P}_{\mathcal{D}_{i}}(\mathsf{X}=\ell_{2})}$, policy $\widehat{\sigma}$ continues to follow policy $\sigma_{\mathbf{{E}}}$’s branch under the outcome $\ell_{1}$ (i.e., it wastes $\ell_{1}-\ell_{2}$ space in the knapsack). Note that the total probability $\sum_{\ell_{1}\geq\ell_{2}}\frac{f_{\ell_{1},\ell_{2}}^{i}}{\mathbf{P}_{\mathcal{D}_{i}}(\mathsf{X}=\ell_{2})}=1$ by definition of the $b$-transportation. Moreover, the probability of reaching any node $s$ in policy $\sigma_{\mathbf{{E}}}$ (under $\mathbf{E}$) equals the probability of reaching the same node $s$ in policy $\widehat{\sigma}$ (under $\mathbf{D}$). Thus, the expected rewards of the two policies are the same, $\mathbb{E}_{\mathbf{{X}}\sim\mathbf{{D}}}[f(\widehat{\sigma},\mathbf{{X}})]=\mathbb{E}_{\mathbf{{X}}\sim\mathbf{{E}}}[f(\sigma_{\mathbf{{E}}},\mathbf{{X}})]$. We note that $\widehat{\sigma}$ is a randomized policy; however, as noted in §1.4, there is always an optimal deterministic policy for these stochastic problems. So, we obtain $\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})\geq\mathtt{OPT}_{\mathcal{I}}(\mathbf{{E}})$ as needed. ∎
Stochastic Matching and Probing
@@
In the stochastic matching problem [${}^{+}$, ${}^{+}$12], there is an undirected graph $G=(V,E)$ with edge-weights $\{w_{e}\}_{e\in E}$, edge-probabilities $\{p_{e}\}_{e\in E}$, and vertex bounds $\{t_{v}\}_{v\in V}$. Each edge $e$ is active independently with probability $e$. However, the status (active/inactive) of an edge can only be determined by probing it. There is also a constraint on the set of probed edges: for any vertex $v$, the number of probed edges incident to $v$ must be at most $t_{v}$. A solution/policy needs to to probe a subset of edges and select a matching $M$ consisting of active edges. The objective is to maximize the expected weight of $M$. Finally, there is a “query commit” requirement that any probed edge which is active must be included in the selected matching $M$. Observe that in any policy execution, we only see the status of some subset of edges, which corresponds to semi-bandit feedback. This problem has been extensively studied, see e.g. [${}^{+}$, ${}^{+}$12, Ada11, BGMS21]. The current best approximation ratio is $0.5$ for the unweighted case [Ada11] and $0.382$ for the weighted case  [BGMS21]. We show in Lemma 3.3 below that the stochastic matching problem is down-monotone. So, Theorem 1.1 with $k=2$ (as all r.v.s are binary) implies an online learning algorithm for stochastic matching having $0.382$-regret of $O(nW\sqrt{T\log T})$ where $W=\sum_{e\in E}w_{e}$ is the total weight.
@@
In fact, our result applies to the much more general stochastic probing problem, as defined in [GN13]. Here, we have a set $E$ of stochastic items with weights $\{w_{e}\}_{e\in E}$ and probabilities $\{p_{e}\}_{e\in E}$. Each item is active independently with probability $e$, and this status can only be determined by probing $e$. We now have two downward-closed constraints: an inner constraint ${\cal F}_{in}$ and an outer constraint ${\cal F}_{out}$. We want to probe a set $Q$ of items subject to the outer constraint (i.e. $Q\in{\cal F}_{out}$) and select a subset $S\subseteq Q$ of active (probed) items satisfying the inner constraint (i.e. $S\in{\cal F}_{in}$). The objective is to maximize the expected weight of the selected items $S$. We again have the query-commit requirement that any active probed item must be selected into the solution $S$. When both inner/outer constraints are $k_{in}$ and $k_{out}$ systems (see [GN13] for the definition of $k$-systems), there is a $\frac{1}{k_{in}+k_{out}}$ approximation for the unweighted case and an $\Omega(\frac{1}{(k_{in}+k_{out})^{2}})$ approximation for the weighted case [GN13]. When the $k$-systems are intersections of matroids, [ASW16] gave an improved adaptive algorithm with approximation ratio $\frac{1}{k_{in}+k_{out}}$, even for the weighted case.
@@
Let $\sigma$ denote the optimal policy for $\mathcal{I}$ under distribution $\mathbf{{D}}$, and let $\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})$ denote its expected weight. Recall that a policy is given by a decision tree, where each node is labeled by an item to probe next, and the branches out of a node correspond to the random $0-1$ realization of the probed item. We construct a policy $\widehat{\sigma}$ that works under distribution $\mathbf{{E}}$, by closely following $\sigma$. When policy $\sigma$ probes item $e$ at any node, policy $\widehat{\sigma}$ does the following:
@@
Note that conditioned on reaching such a node, the probability that $\widehat{\sigma}$ goes to the active branch is $\frac{p_{e}}{\bar{p}_{e}}\cdot\bar{p}_{e}=p_{e}$; so the probability that $\widehat{\sigma}$ goes to the inactive branch is $1-p_{e}$. Therefore, the probability of reaching any node $v$ in policy $\widehat{\sigma}$ (under distribution $\mathbf{E}$) equals the probability of reaching node $v$ in policy ${\sigma}$ (under distribution $\mathbf{D}$). Thus the expected values of these two policies are the same, $\mathbb{E}_{\mathbf{{x}}\sim\mathbf{{E}}}[f(\widehat{\sigma},\mathbf{{x}})]=\mathbb{E}_{\mathbf{{x}}\sim\mathbf{{D}}}[f(\sigma,\mathbf{{x}})]$. We note that $\widehat{\sigma}$ also satisfies the query-commit requirement because $\sigma$ does (whenever $e$ is probed and found active, it will be included in the solution). Finally, observe that $\widehat{\sigma}$ is a randomized policy; however, as noted in §1.4, there is always an optimal deterministic policy for these stochastic problems. So, we obtain $\mathtt{OPT}_{\mathcal{I}}(\mathbf{{E}})\geq\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})$ as needed. ∎
Stochastic Covering Knapsack
Here, we have $n$ items with deterministic costs $\{c_{i}\}_{i\in[n]}$ and (independent) random rewards $\{\mathsf{R}_{i}\}_{i\in[n]}$. There is a known probability distribution $\mathcal{D}_{i}$ for $\mathsf{R}_{i}$; we assume discrete distributions here. Given a target $Q$, we need to select items until the total observed reward is at least $Q$ (or we run out of items). The objective is to minimize the expected cost. Note that we only observe items until the target is achieved, which corresponds to semi-bandit feedback. There is an adaptive $3$-approximation algorithm for this problem [DHK16]. By a proof identical to that of Lemma 3.1, we get:
@@
So, using Theorem 1.1, we obtain a polynomial time online learning algorithm for stochastic min-knapsack having $3$-regret $O(nkC\sqrt{T\log(kT)})$ where $C=\sum_{i=1}^{n}c_{i}$ is the total cost and $k$ is the maximum support size.
Stochastic $k$-TSP
@@
The stochastic $k$-TSP problem is the stochastic variant of the classical $k$-TSP problem [Gar05]. An instance of the problem comprises a metric $(V,d)$ with a root vertex $r\in V$. There are random rewards $\{\mathsf{R}_{v}\}_{v\in V}$ associated with the vertices. All reward distributions are given as input, but the realized reward $\mathsf{R}_{v}$ is only known when vertex $v$ is visited. We assume that all reward distributions are discrete. Given a target $Q$, the goal is to find an (adaptive) tour originating from $r$ that collects a total reward at least $Q$ (or runs out of vertices), at the minimum expected length. A constant-factor approximation algorithm for stochastic $k$-TSP was obtained in [JLLS20], which was via a non-adaptive policy (so it is 1-threshold). Moreover, using a proof identical to that of Lemma 3.4, it follows that this problem is also up-monotone. Now, using Theorem 2.9, we obtain a polynomial time online learning algorithm for stochastic $k$-TSP having $O(1)$-regret $O(nD\sqrt{T\log T})$ where $D=\sum_{v\in V}d_{rv}$ corresponds to an upper bound on the (full) TSP cost.
Stochastic Submodular Maximization.
Consider a monotone submodular function $f:2^{E}\rightarrow\mathbb{Z}_{+}$ defined on a groundset $E$. Constrained submodular maximization involves selecting a subset $S\subseteq E$ of items in a downward-closed family ${\cal F}\subseteq 2^{E}$, that maximizes the function value $f(S)$. This is a fundamental problem in combinatorial optimization and good approximation algorithms are known for a variety of constraints ${\cal F}$ such as cardinality, matroids and knapsacks [NWF78, CCPV11, KST13, CVZ14]. We consider a stochastic variant, as in [AN16, GNS17, BSZ19], where each item $i\in E$ is active independently with some probability $p_{i}$ (and inactive otherwise). [^5] These papers also handle a more general setting where items have an arbitrary outcome space (not just binary). Moreover, the function value is accrued only on active items. The active/inactive status of any item is only known after selecting it. The goal now is to select a subset $S\in{\cal F}$ sequentially (and adaptively) so as to maximize the expected function value from the selected active items, i.e., $\mathbb{E}_{S,A}\left[f(S\cap A)\right]$ where $A\subseteq E$ is the random set of active items. Note that we can only select (and observe) a constrained subset of items, which corresponds to semi-bandit feedback. When ${\cal F}$ is a single matroid constraint, [AN16] obtained a $(1-\frac{1}{e}-\epsilon)$-approximation algorithm for the stochastic submodular maximization problem. Later, [GNS17, BSZ19] showed that a $\rho$-approximation algorithm for deterministic submodular maximization under any constraint ${\cal F}$ can be used to obtain a $\frac{\rho}{2}$-approximation algorithm for the stochastic problem under constraint ${\cal F}$. It is easy to see that this problem is down-monotone. So, Theorem 1.1 with $k=2$ (binary r.v.s) implies a $(1-\frac{1}{e}-\epsilon)$-regret of $O(nF\sqrt{T\log T})$ for stochastic submodular maximization over a matroid constraint, where $F$ is the maximum function value. Combined with the result of [BSZ19], we also obtain online learning algorithms for more general constraints ${\cal F}$.
Stochastic Submodular Cover

### v2

@@ 5.4 Stochastic Knapsack
The proof of this lemma requires the following theorem (also referred to as Strassen’s monotone coupling theorem) on coupling random variables under stochastic dominance (Strassen, 1965; Liggett and Liggett, 1985).
Theorem 5.2 (Coupling under Stochastic Dominance).
@@
Consider r.v.s $\mathsf{X}$ and $\mathsf{Y}$ with finite support $k]=\{1,\ldots,k\}$, and distributed according to $\mathcal{D}$ and $\mathcal{E}$ respectively. Furthermore, suppose that $\mathcal{D}\preceq_{\mathsf{SD}}\mathcal{E}$. Then, there exist values $\left\{f_{\ell_{1},\ell_{2}}:k\geq\ell_{1}\geq\ell_{2}\geq 1\right\}$ such that:
$\sum_{\ell_{2}:\ell_{2}\leq\ell_{1}}f_{\ell_{1},\ell_{2}}=\Pr_{\mathcal{E}}(\mathsf{Y}=\ell_{1}),\,\forall\ell_{1}\in[k]\quad\mbox{ and }\quad\sum_{\ell_{1}:\ell_{1}\geq\ell_{2}}f_{\ell_{1},\ell_{2}}=\Pr_{\mathcal{D}}(\mathsf{X}=\ell_{2}),\,\forall\ell_{2}\in[k].$
Proof of Lemma 5.1.
@@
Consider an arbitrary instance $\mathcal{I}$ and product probability distributions $\mathbf{{D}}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ and $\mathbf{{E}}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ where $\mathbf{{D}}\preceq_{\mathsf{SD}}\mathbf{{E}}$. For each item $i\in[n]$, let $f_{\ell_{1},\ell_{2}}^{i}$ denote coupled values for distributions $\mathcal{D}_{i}\preceq_{\mathsf{SD}}\mathcal{E}_{i}$, constructed as per Theorem 5.2. Let $\sigma_{\mathbf{{E}}}$ denote the optimal policy for $\mathcal{I}$ under distribution $\mathbf{{E}}$, and let $\mathtt{OPT}_{\mathcal{I}}(\mathbf{{E}})$ denote its reward. Recall that a policy is given by a decision tree, where each node is labeled by an item to probe next, and the branches out of a node correspond to the random realization of the probed item. We construct a policy $\widehat{\sigma}$ that works under distribution $\mathbf{{D}}$, by closely following $\sigma_{\mathbf{{E}}}$. When policy $\sigma_{\mathbf{{E}}}$ probes some item $i$, policy $\widehat{\sigma}$ does the same. Now, suppose that $\widehat{\sigma}$ observes $C_{i}=\ell_{2}\in[k]$. Then, for each $\ell_{1}\geq\ell_{2}$, with probability $\frac{f_{\ell_{1},\ell_{2}}^{i}}{\Pr_{\mathcal{D}_{i}}(\mathsf{X}=\ell_{2})}$, policy $\widehat{\sigma}$ continues to follow policy $\sigma_{\mathbf{{E}}}$’s branch under the outcome $\ell_{1}$ (i.e., it wastes $\ell_{1}-\ell_{2}$ space in the knapsack). Note that the total probability $\sum_{\ell_{1}\geq\ell_{2}}\frac{f_{\ell_{1},\ell_{2}}^{i}}{\Pr_{\mathcal{D}_{i}}(\mathsf{X}=\ell_{2})}=1$ by definition of the $b$-transportation. Moreover, the probability of reaching any node $s$ in policy $\sigma_{\mathbf{{E}}}$ (under $\mathbf{E}$) equals the probability of reaching the same node $s$ in policy $\widehat{\sigma}$ (under $\mathbf{D}$). Thus, the expected rewards of the two policies are the same, $\mathbb{E}_{\mathbf{{X}}\sim\mathbf{{D}}}[f(\widehat{\sigma},\mathbf{{X}})]=\mathbb{E}_{\mathbf{{X}}\sim\mathbf{{E}}}[f(\sigma_{\mathbf{{E}}},\mathbf{{X}})]$. We note that $\widehat{\sigma}$ is a randomized policy; however, as noted in §1.5, there is always an optimal deterministic policy for these stochastic problems. So, we obtain $\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})\geq\mathtt{OPT}_{\mathcal{I}}(\mathbf{{E}})$ as needed. ∎
5.5 Stochastic Matching and Probing
@@
In the stochastic matching problem (Chen et al., 2009; Bansal et al., 2012), there is an undirected graph $G=(V,E)$ with edge-weights $\{w_{e}\}_{e\in E}$, edge-probabilities $\{p_{e}\}_{e\in E}$, and vertex bounds $\{t_{v}\}_{v\in V}$. Each edge $e$ is active independently with probability $e$. However, the status (active/inactive) of an edge can only be determined by probing it. There is also a constraint on the set of probed edges: for any vertex $v$, the number of probed edges incident to $v$ must be at most $t_{v}$. A solution/policy needs to to probe a subset of edges and select a matching $M$ consisting of active edges. The objective is to maximize the expected weight of $M$. Finally, there is a “query commit” requirement that any probed edge which is active must be included in the selected matching $M$. Observe that in any policy execution, we only see the status of some subset of edges, which corresponds to semi-bandit feedback. This problem has been extensively studied, see e.g., Chen et al. (2009); Bansal et al. (2012); Adamczyk (2011); Brubach et al. (2021). The current best approximation ratio is $0.5$ for the unweighted case (Adamczyk, 2011) and $0.382$ for the weighted case (Brubach et al., 2021). We show in Lemma 5.3 below that the stochastic matching problem is down-monotone. So, Theorem 1.1 with $k=2$ (as all r.v.s are binary) implies an online learning algorithm for stochastic matching having $0.382$-regret of $O(nW\sqrt{T\log T})$ where $W=\sum_{e\in E}w_{e}$ is the total weight.
@@
In fact, our result applies to the much more general stochastic probing problem, as defined in Gupta and Nagarajan (2013). Here, we have a set $E$ of stochastic items with weights $\{w_{e}\}_{e\in E}$ and probabilities $\{p_{e}\}_{e\in E}$. Each item is active independently with probability $p_{e}$, and this status can only be determined by probing $e$. We now have two downward-closed constraints: an inner constraint ${\cal F}_{in}$ and an outer constraint ${\cal F}_{out}$. We want to probe a set $Q$ of items subject to the outer constraint (i.e. $Q\in{\cal F}_{out}$) and select a subset $S\subseteq Q$ of active (probed) items satisfying the inner constraint (i.e. $S\in{\cal F}_{in}$). The objective is to maximize the expected weight of the selected items $S$. We again have the query-commit requirement that any active probed item must be selected into the solution $S$. When both inner/outer constraints are $k_{in}$ and $k_{out}$ systems, there is a $\frac{1}{k_{in}+k_{out}}$ approximation for the unweighted case and an $\Omega(\frac{1}{(k_{in}+k_{out})^{2}})$ approximation for the weighted case (Gupta and Nagarajan, 2013). When the $k$-systems are intersections of matroids, Adamczyk et al. (2016) gave an improved adaptive algorithm with approximation ratio $\frac{1}{k_{in}+k_{out}}$, even for the weighted case.
@@
Let $\sigma$ denote the optimal policy for $\mathcal{I}$ under distribution $\mathbf{{D}}$, and let $\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})$ denote its expected weight. Recall that a policy is given by a decision tree, where each node is labeled by an item to probe next, and the branches out of a node correspond to the random 010-1 realization of the probed item. We construct a policy $\widehat{\sigma}$ that works under distribution $\mathbf{{E}}$, by closely following $\sigma$. When policy $\sigma$ probes item $e$ at any node, policy $\widehat{\sigma}$ does the following:
@@
Note that conditioned on reaching such a node, the probability that $\widehat{\sigma}$ goes to the active branch is $\frac{p_{e}}{\bar{p}_{e}}\cdot\bar{p}_{e}=p_{e}$; so the probability that $\widehat{\sigma}$ goes to the inactive branch is $1-p_{e}$. Therefore, the probability of reaching any node $v$ in policy $\widehat{\sigma}$ (under distribution $\mathbf{E}$) equals the probability of reaching node $v$ in policy $\sigma$} (under distribution $\mathbf{D}$). Thus the expected values of these two policies are the same, $\mathbb{E}_{\mathbf{{x}}\sim\mathbf{{E}}}[f(\widehat{\sigma},\mathbf{{x}})]=\mathbb{E}_{\mathbf{{x}}\sim\mathbf{{D}}}[f(\sigma,\mathbf{{x}})]$. We note that $\widehat{\sigma}$ also satisfies the query-commit requirement because $\sigma$ does (whenever $e$ is probed and found active, it will be included in the solution). Finally, observe that $\widehat{\sigma}$ is a randomized policy; however, as noted in Section 1.5, there is always an optimal deterministic policy for these stochastic problems. So, we obtain $\mathtt{OPT}_{\mathcal{I}}(\mathbf{{E}})\geq\mathtt{OPT}_{\mathcal{I}}(\mathbf{{D}})$ as needed. ∎
5.6 Stochastic Covering Knapsack
Here, we have $n$ items with deterministic costs $\{c_{i}\}_{i\in[n]}$ and (independent) random rewards $\{\mathsf{R}_{i}\}_{i\in[n]}$. There is a known probability distribution $\mathcal{D}_{i}$ for $\mathsf{R}_{i}$; we assume discrete distributions here. Given a target $Q$, we need to select items until the total observed reward is at least $Q$ (or we run out of items). The objective is to minimize the expected cost. Note that we only observe items until the target is achieved, which corresponds to semi-bandit feedback. There is an adaptive $3$-approximation algorithm for this problem (Deshpande et al., 2016). By a proof identical to that of Lemma 5.1, we get:
@@ 5.6 Stochastic Covering Knapsack
So, using Theorem 1.1, we obtain a polynomial time online learning algorithm for stochastic min-knapsack having $3$-regret $O(nkC\sqrt{T\log T})$ where $C=\sum_{i=1}^{n}c_{i}$ is the total cost and $k$ is the maximum support size.
Stochastic $\kappa$-TSP.
@@
This is the stochastic variant of the classical $\kappa$-TSP problem (Garg, 2005). An instance of the problem comprises a metric $(V,d)$ with a root vertex $r\in V$. There are random rewards $\{\mathsf{R}_{v}\}_{v\in V}$ associated with the vertices. All reward distributions are given as input, but the realized reward $\mathsf{R}_{v}$ is only known when vertex $v$ is visited. We assume that all reward distributions are discrete. Given a target $Q$, the goal is to find an (adaptive) tour originating from $r$ that collects a total reward at least $Q$ (or runs out of vertices), at the minimum expected length. A constant-factor approximation algorithm for stochastic $\kappa$-TSP was obtained in Jiang et al. (2020). Applying Theorem 1.1, we obtain a polynomial time online learning algorithm for stochastic $\kappa$-TSP having $O(1)$-regret $O(nkD\sqrt{T\log T})$ where $D=\sum_{v\in V}d_{rv}$ corresponds to an upper bound on the (full) TSP cost.
5.7 Stochastic Submodular Optimization
[…]
Stochastic Submodular Cover.